Sin s= 3/5 sin t= 12/13 quadrant third
Found 2 solutions by MathLover1, ikleyn:
Answer by MathLover1(20850)


question is incomplete
what is to find?

Answer by ikleyn(52803)

.
Sin s= 3/5 sin t= 12/13 quadrant third
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


For your info,  dear problem's  creator:

        in third quadrant,  sine function is always  NEGATIVE  and can not have positive values.

        So,  the data in your post is  DEFECTIVE.

        Learn the basics of  Trigonometry.